Let F(x) = Sý sin*(0) dt. Evaluate the following limit . Let F(x) = $* sin?(t) dt. Evaluate the following limit. 2022
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

a lim xso F(x) 2 int dt - x10 Jo lim x2 xso - So sink it dt = o form 10i F(x) = dimr a rra 10x2 xrol de Jo 34 - Sentusu) 2 (t) det 0 2x because a (x) = nxn- Now by Leibnitz sale: - & pace nicado8 fos e sindt dt] f(x)= 0 g(x)=x ah(t) = sind (t) - 3 Now h[g(x)] = h(x) [because g(x)= x =) h[glx)] = sin? (x) - [by putting(because sin(o)=o] lind Flet - o form again using Dili Hospitals sule: - - lingo Flx) = lim (dhe [sin (x)) d (2x) Vm x70 dX- lim Um xto de (sin(x)) (2x) J da dim e simteshkasine)] because she (sin (x) = n simma (eti ( sinta a (nx) =n because sin(x